OpenStudy (anonymous):

PLEASE HELP! COMPLETE THE SQUARE! x^2-7x=0

OpenStudy (kinggeorge):

Recall that if you have a polynomial of the form\[x^2+bx+c\]then when you complete the square you add and subtract the value \(b^2/4\). In this case, can you tell me what your value of \(b\) is?

OpenStudy (anonymous):

-7

OpenStudy (kinggeorge):

Right. So to complete the square we would have to add\[\frac{(-7)^2}{4}=\frac{49}{4}.\]This results in\[x^2-7x+\frac{49}{4}-\frac{49}{4}=0\]or equivalently\[x^2-7x+\frac{49}{4}=\frac{49}{4}.\]

OpenStudy (anonymous):

i got that so far

OpenStudy (kinggeorge):

Good. Now you have to remember/notice that\[x^2+bx+\frac{b^2}{4}=\left(x+\frac{b}{2}\right)^2.\]Applying that to your problem, since \(b=-7\), we get\[x^2-7x+\frac{49}{4}=\left(x-\frac{7}{2}\right)^2=\frac{49}{4}.\]Still following?

OpenStudy (anonymous):

yup!

OpenStudy (kinggeorge):

Excellent. So now we can take the square root of both sides. Thus,\[x-\frac{7}{2}=\pm\sqrt{49/4}=\pm\frac72.\]This leads to two possibilities:\[x-\frac{7}{2}=\frac{7}{2}\implies x=\frac{7}{2}+\frac72=7\]or\[x-\frac72=-\frac72\implies x=0.\]So your two roots will be \(x=7\) and \(x=0\).

OpenStudy (anonymous):

shoot i wish i couldve done that! can u help me with a few more? ur smart

OpenStudy (kinggeorge):

I've got a bit of time to help out. Make sure you've tried the problems on your own first though!

OpenStudy (anonymous):

oh i have! these are problems i was unable to solve

OpenStudy (anonymous):

the height y of a parabolic arch is given by y=-1/16x^2+40, where x is the horizonal distance from the center of the base of the arch. All distances are in feet! A.) what is the highest point of the arch? B.) How wide is the arch at the base to the nearest tenth of a foot?

OpenStudy (anonymous):

i have no idea how to do this

OpenStudy (kinggeorge):

First you need to remember what the formula is for the center of a parabola. If your equation is\[ax^2+bx+c,\]then the center is given by \(x=-b/2a\). In your case, \(b=0\), so that the center will be at \(x=0\). So we're halfway to finding the maximum height. To find the max height, we need to plug this value back into the equation. So\[y(0)=\frac{-1}{16}0^2+40=40 \;ft.\]is the maximum height. Make sense so far?

OpenStudy (anonymous):

poo kiddo... completing the square on this problem, Lmao .

OpenStudy (anonymous):

yes so far yes.

OpenStudy (kinggeorge):

Good. To find the width at the base, you have to realize that the base is where the height is 0. So you need to find the roots of your equation. Can you do that?

OpenStudy (anonymous):

hmm im not sure...im trying.. so the highest point is 40 but the hieght is still 0?

OpenStudy (kinggeorge):

The highest point is 40, but that was part a. Now we're focusing on the second part. And kind of by definition of what an arch is, the base of the arch should be at height 0. It will meet the ground at two points, say \(r_1\) and \(r_2\). Then the distance between these points will merely be \(r_1-r_2\) (or possibly \(r_2-r_1\) if the plus/minus signs don't work out). So to find the points where the arch meets the ground, we need to set height equal to 0.

OpenStudy (anonymous):

ive been trying all kinds of ways to get this...ugh i suck at math

OpenStudy (anonymous):

ok ok ..so how do we do that? put it into our formula?

OpenStudy (kinggeorge):

What formula? The quadratic formula?

OpenStudy (anonymous):

i guess ya

OpenStudy (kinggeorge):

That would work just fine (although is a slightly faster way in this case). Can you tell me what you get for the roots?

OpenStudy (anonymous):

huh? u have to show me i dont understand

OpenStudy (anonymous):

like where would u put the 40?

OpenStudy (kinggeorge):

So your equation is \[-\frac{x^2}{16}+40\]To find the roots we set it equal to 0\[-\frac{x^2}{16}+40=0\]So\[40=\frac{x^2}{16}\]Take the square root of both sides, and\[\pm\sqrt{40}=\pm2\sqrt{10}=\frac{x}{4}\]So\[x=\pm8\sqrt{10}\]Did you follow all this?

OpenStudy (anonymous):

the answer is suppsed to be 50.6 but idk how the answer key got it

OpenStudy (kinggeorge):

We're almost there. Now we find the distance between those two values. The distance is just\[8\sqrt{10}-(-8\sqrt{10})=8\sqrt{10}+8\sqrt{10}=16\sqrt{10}.\]When you use a calculator to find this value numerically, we get\[16\sqrt{10}\approx50.6\]

OpenStudy (anonymous):

see i did this all wrong, i did it quadratic formula way and ya it didnt work

OpenStudy (kinggeorge):

The quadratic formula way should still work, but there's more chance for error.

OpenStudy (anonymous):

i think i probably did then, i like your way better.my teacher didnt teach me a good way ):

OpenStudy (kinggeorge):

That's too bad... A lot of this just comes with practice, experience, and seeing other people do the same thing before me though.

OpenStudy (anonymous):

oh really? hmm. i just need more practice. Do you know how to solve this? (8-3i)(6+9i)??? you Foil right?

OpenStudy (kinggeorge):

Right. Just remember that \(i^2=-1\).

OpenStudy (anonymous):

i got 75+54i?

OpenStudy (kinggeorge):

Looks right.

OpenStudy (anonymous):

how do you solve y=x^2-11x+24 y=x-3 Solve each system

OpenStudy (kinggeorge):

Why don't you post this in a new thread now so that we can start with a blank slate again.

OpenStudy (anonymous):

ok